One of two numbers is five more than the other. the sum of the numbers is 17. find the numbers
Answer by nanobot101(2)

let
x be the larger number and y be the smaller one.
x+5=y and x+y=17
use substitution:
x+(5+x)=17
2x+5=17
2x=12
x=6
sub x=6 into the first equation.
(6)+5=y
11=y. there fore the solution is (6,11)